OpenStudy (asib1214):

ok if you overheat the CuSO4*5H2O compound and all of the water molecules escapes/evaporate. After that you apply heat further more to the remaining compound CuSO4 (anhydrous) so that about 10% of the Copper becomes Cu(II) oxide. two questions here, what is the ratio of water to CuSO4 (after evaporation) and what would be the chemical equation of this reaction??? at the end???

OpenStudy (asib1214):

the total mass of the hydrate is 11.28g, in which about 7.28g is CUSO4 and 4.1g is water.

OpenStudy (unklerhaukus):

can you convert those masses to moles?

OpenStudy (asib1214):

i've already done that the ratio i get is 1to5

OpenStudy (asib1214):

another way to get the ratio is if you divide 4 by 7.28....

OpenStudy (asib1214):

4.1

OpenStudy (asib1214):

i just wanna know what is the chemical equation of this compound???? is it CuSO4 + O2 = CUO + SO

OpenStudy (unklerhaukus):

wouldent it be sulfur dioxide?

OpenStudy (asib1214):

CuSO4= 7.28/159.62=0.456 moles H2O = 4.1/18=0.227 moles

OpenStudy (asib1214):

i know for sure that the copper will become Cu(II) oxide and yah maybe the remaining sulfur will become SO2...

OpenStudy (asib1214):

because it says in the question...that the hydrate will produce "some" copper(II) oxide.

OpenStudy (unklerhaukus):

i think you'll get CuSO4 + O2 = CuO + SO2

OpenStudy (asib1214):

thats what i have too bah i'm not too sure :/ aight i'll take that as a right answer :)

OpenStudy (asib1214):

and what about the ratio of water to anhydrous????

OpenStudy (asib1214):

CuSO4